With a height of 69 ​in, William was the shortest president of a particular club in the past century. The club presidents of the past century have a mean height of 74.8 in and a standard deviation of 1.9 in.
a.What is the positive difference between William​'s height and the​ mean?
b.How many standard deviations is that​ [the difference found in part​ (a)]?
c.Convert William​'s height to a z score.
d.If we consider​ "usual" heights to be those that convert to z scores between -2 and​ 2, is William​'s height usual or​ unusual?
A. The positive difference between William​'s height and the mean is ?
5.8 in.
B. The difference is__standard deviations?
​(Type an integer or decimal rounded to two decimal places as​ needed.)
c. The z score is___?
​(Type an integer or decimal rounded to two decimal places as​ needed.)
D. Is William​'s height usual or​ unusual?

Answer by stanbon(75887)

With a height of 69 ​in, William was the shortest president of a particular club in the past century. The club presidents of the past century have a mean height of 74.8 in and a standard deviation of 1.9 in.
a.What is the positive difference between William​'s height and the​ mean?
74.8-69 = 5.8"
----------------------------
b.How many standard deviations is that​ [the difference found in part​ (a)]?
5.8/1.9 = 3.05
----------------------
c.Convert William​'s height to a z score.
z(69) = (69-74.8)/1.9 = -3.05
-----------------------------------
d.If we consider​ "usual" heights to be those that convert to z scores between -2 and​ 2, is William​'s height usual or​ unusual?
Ans: unusual because -3.05 is outside that range.
-----------------------
Cheers,
Stan H.
-------------